OpenStudy (anonymous):

can someone please help me with a simple interest problem? A woman invested $9000 in two mutual funds. In one year part of her investment earned 5% simple interest and the other part earned 6% simple interest. At the end of that year, the woman had earned $471. How much did she invest at each rate?

OpenStudy (dan815):

x+y=9000 0.05*x+0.06*y=471

OpenStudy (dan815):

you must solve these system of equations

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

i just never understand how to set it up

OpenStudy (anonymous):

thankyou kind sir

OpenStudy (dan815):

You are welcome :)

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

i have those equations, but I'm not sure what to do with them now...

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

Okay so, we can solve for Y in terms of x, in the first equation and plug that into the 2nd equation and solve for x

OpenStudy (anonymous):

and how do you do that

OpenStudy (dan815):

\[x+y=9000\\ y=9000-x\\ 0.05*x+0.06*y=471\\ 0.05*x+0.06*(9000-x)=471\\ -0.01x+0.06(9000)=471\\ x=\frac{(471-0.06*9000)}{-0.01}\\ after ~you~find~'x'~you~ can~ find~'y'~by~x+y=9000~~equation \]